OpenStudy (ny,ny):

Find all points of inflection of the graph of the function f(x) = (1/12)x^4-2x^2+15

OpenStudy (ny,ny):

I found the first derivative to be (1/3)x^3-4x And the second derivative to be x^2-4 And when f"(x) = 0, then x=2,-2

OpenStudy (mathmale):

1) Find the first and second derivatives, f '(x) and f "(x). 2) Set f "(x) = to 0. 3) Solve the resulting equation. In this case you'll obtain 2 roots. 4) Determine whether the second derivative changes sign at either or both of these 2 roots. If yes, you have an inflection point there; if no, you don't. Pls show your work.

OpenStudy (ny,ny):

"Determine whether the second derivative changes sign at either or both of these 2 roots." You mean plug in a number less than -2, between -2 and 2, and greater than 2 to see if theres a change in sign between these three values, right?

OpenStudy (ny,ny):

And if so, plug it into the original equation?

OpenStudy (karim728):

whats the second derivative?

OpenStudy (ny,ny):

I said that the second derivative is x^2-4 @mathmale @karim728

OpenStudy (doodleduck8):

Super easy...all you need is the 2nd derivative!! f ' (x) = (1/3)x^3 - 4x f '' (x) = x^2 - 4 now set the 2nd derivative equal to 0 and solve for x: f '' (x) = 0 = x^2 - 4 x^2 = 4 x = + or - 2 You have two points of inflection: ( -2, f(-2) ) and ( 2, f(2) ). Be sure to substitute the value for x into the original equation to find your y.

OpenStudy (ny,ny):

Oh ok. So I got (-2,25/3) and (2,25/3) Oh thank you so much for your help, I was waiting so long.

OpenStudy (doodleduck8):

No problem!! :)